OpenStudy (shesolitt):

I WILL MEDAL AND FAN Find the average rate of change for the given function from x = −1 to x = 2. Graph of f of x equals negative three x plus 2 negative one-third one-third −3 3

OpenStudy (will.h):

Post the graph

OpenStudy (will.h):

Hint rate of change can be found using Y2-Y1/X2-X1

OpenStudy (will.h):

F(x) = -3x + 2 That's the function They want to know when x = -1 and when x = 2 First find the y value of each x value Start with x = -1 -3(-1) + 2 Y = 5 Now when x = 2 -3(2) + 2 Y = -4 So we have (-1,5) (2,-4) Average rate of change = Y2-Y1/X2-X1 Average = -4 -5/2-(-1) Average = -9/3 Average = -3 Make sense?

OpenStudy (shesolitt):

YES THANK YOU OMGGG

OpenStudy (will.h):

You welcome :)